dx/dt=x-y dy/dt=x +3y

1
Expert's answer
2021-05-03T07:18:46-0400
"\dfrac{dx}{dt}=x-y"

"\dfrac{dy}{dt}=x+3y"

"\begin{pmatrix}
 dx/dt \\
 dy/dt
\end{pmatrix}=\begin{pmatrix}
 1 & -1 \\
 1 & 3
\end{pmatrix}
\begin{pmatrix}
 x \\
 y
\end{pmatrix}"

Find the eigenvalues 


"\begin{pmatrix}
 1-\lambda & -1 \\
 1 & 3-\lambda
\end{pmatrix}=0"

"(1-\lambda)(3-\lambda)+1=0"

"\lambda^2-4\lambda+4=0"

"\lambda_1=\lambda_2=2"

 Find the eigenvectors.

"\lambda=2"


"\begin{pmatrix}
 1-2 & -1 \\
 1 & 3-2
\end{pmatrix}\begin{pmatrix}
 v_1 \\
 v_2
\end{pmatrix}=0"

The eigen vector is


"v=\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}"

For repeated eigenvalues "\lambda_1=\lambda_2=2" and the eigenvector "v=\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}"

the general solution takes form


"\begin{pmatrix}
 x \\
 y
\end{pmatrix}=c_1e^{2t}\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}+c_2 te^{2t}\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}+c_2e^{2t}\begin{pmatrix}
 u_1 \\
 u_2
\end{pmatrix}"

where


"\begin{pmatrix}
 1-2 & -1 \\
 1 & 3-2
\end{pmatrix}\begin{pmatrix}
 u_1 \\
 u_2
\end{pmatrix}=\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}"

"\begin{pmatrix}
 u_1 \\
 u_2
\end{pmatrix}=\begin{pmatrix}
 1 \\
 0
\end{pmatrix}"

The solution is


"\begin{pmatrix}
 x \\
 y
\end{pmatrix}=c_1e^{2t}\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}+c_2 te^{2t}\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}+c_2e^{2t}\begin{pmatrix}
 1 \\
 0
\end{pmatrix}"